DIRC Operations Weekly Report
Period: Wednesday 2007/6/19 - Wednesday 2007/6/27
Oncall: Justine Serrano
Fast Monitoring: Justine Serrano, Alejandro Perez
Luminosities this week
PEP peak lumi: ~1 x 1034
(last week: ~1 x 1034)
PEP delivery: 1.92 fb-1
(last weeks: 2.27 fb-1)
BaBar integrated lumi:1.76 fb-1
(last weeks: 2.2 fb-1)
Run 6 total delivery: 51.02 fb-1
Run 6 total BaBar recorded: 49.05 fb-1 total, including 4.18 fb-1 off-peak data
PEP activities/issues this week
06/20: transition from OffPeak -> OnPeak.
06/21 to 06/22: ROD. They replace a bellow in arc3. Leak discovered in sec 20 of the Linac, they decided not to fix it during the ROD. Starting injecting on 06/22 at 17:05.
06/23: PEP recovered a normal luminosity.
BaBar activities/issues this week
06/21: ROD. Tried to fix a EH cooling problem but the spare didn't fit.
06/24: DCH had a problem with an electronic box inside the DIRC tunnel. We need a new ROD to fix that.
DIRC activities/issues this week
06/21: ROD. We turned off the HV, FEE crates and the chiller. Mike replaced a connector card at the back of the chiller power siam.
06/21, while BaBar was starting to take cosmics, we lost the monitoring of sector 5. Calibration and a cosmic run showed that data is still ok. The corresponding alarm has been masked in epics and in the alarm handler, and DIRC has been made runnable again by bypassing sector 5 monitoring.
After the ROD, HV group 1 in sector 0 tripped because of the Xmas tree. Jose set the runnable value at 900V so that it didn't trip anymore.
See DIRC logbook for details
Lumi and Background this week
- ROD scheduled for June 27 (~ 1.5 shift long)
- We will replace the DCC in sector 5. Spare to be used: 26.
- Replace DFB 1 in sector 6. Spares that can be used: 3D
- Unplug X-Tree in sector 0
- Unplug X-Tree in sector 7
- Replace four DFBs with goups of 8 dead or flaky channels:
- sector 0, DFB 0.
- 2 in sector 1, DFB 1 and DFB 13.
- sector 3, DFB 6.
- sector 8, DFB 0.
Alejandro, Justine and Jose rotating backup & on-call commissioning duties.